### v2

# Packing Input Frame Context in Next-Frame Prediction Models for Video Generation

Abstract
We present a neural network structure, FramePack, to train next-frame (or next-frame-section) prediction models for video generation. The FramePack compresses input frames to make the transformer context length a fixed number regardless of the video length. As a result, we are able to process a large number of frames using video diffusion with computation bottleneck similar to image diffusion. This also makes the training video batch sizes significantly higher (batch sizes become comparable to image diffusion training). We also propose an anti-drifting sampling method that generates frames in inverted temporal order with early-established endpoints to avoid exposure bias (error accumulation over iterations). Finally, we show that existing video diffusion models can be finetuned with FramePack, and their visual quality may be improved because the next-frame prediction supports more balanced diffusion schedulers with less extreme flow shift timesteps.
1 Introduction
Forgetting and drifting are the two most critical problems in next-frame (or next-frame-section) prediction models for video generation. Herein, "forgetting" refers to the fading of memory as the model struggles to remember earlier content and maintain consistent temporal dependencies, whereas "drifting" refers to the iterative degradation of visual quality due to error accumulation over time (also called exposure bias).
A fundamental dilemma emerges when attempting to simultaneously address both forgetting and drifting: any method that mitigates forgetting by enhancing memory may also make error accumulation/propagation faster, thereby exacerbating drifting; any method that reduces drifting by interrupting error propagation and weakening the temporal dependencies (e.g., masking or re-noising the history) may also worsen the forgetting. This essential trade-off hinders the scalability of next-frame prediction models.
The forgetting problem leads to a naive solution to encode more frames, but this quickly becomes computationally intractable due to the quadratic attention complexity of transformers (or sub-quadratic optimizations like FlashAttn, etc.). Moreover, video frames contain significant temporal redundancy, rendering naive full-context approaches less efficient. The substantial duplication of visual features across consecutive frames reveals the potential to design effective compression systems to facilitate memorization.
The drifting problem is influenced by memorizing mechanisms from multiple aspects. The source of drifting lies in initial errors that occur in individual frames, while the effect is the propagation and accumulation of these errors across subsequent frames, leading to degraded visual quality. On one hand, a stronger memorizing mechanism leads to better temporal consistency and reduces the occurrence of initial errors, thereby mitigating drifting. On the other hand, a stronger memorizing mechanism also memorizes more errors and thus accelerates error propagation when errors do occur, exacerbating drifting. This paradoxical relationship between memory mechanisms and drifting necessitates carefully designed training/sampling methods to facilitate error correction or interrupt error propagation.
In this paper, we propose FramePack as an anti-forgetting memory structure along with anti-drifting sampling methods. The FramePack structure addresses the forgetting problem by compressing input frames based on their relative importance, ensuring the total transformer context length converges to a fixed upper bound regardless of video duration. This enables the model to encode significantly more frames without increasing the computational bottleneck, thereby facilitating anti-forgetting. Moreover, we propose anti-drifting sampling methods that break the causal prediction chain and incorporate bi-directional context. These methods include generating endpoint frames before filling intermediate content, and an inverted temporal sampling approach where frames are generated in reverse order, each attempting to approach a known high-quality frame (e.g., an input frame in image-to-video tasks). We show that these methods effectively reduce the occurrence of errors and prevent their propagation.
We demonstrate that existing pretrained video diffusion models (e.g., HunyuanVideo, Wan, etc.) can be finetuned with FramePack. Our experiments reveal several findings: because next-frame prediction generates smaller tensor sizes per step compared to full-video generation, it enables more balanced diffusion schedulers with less extreme flow shift timesteps. We observe that the resulting less aggressive schedulers may lead to improved visual quality, beyond the direct goals of addressing the forgetting and drifting problems.
2 Related Work
2.1 Anti-forgetting and Anti-drifting
Noise scheduling and augmentation in history frames modify noise levels at specific timesteps, video times, or image frequencies to create causal computation or anti-drifting effects. These methods generally reduce dependency on past frames. DiffusionForcing [4] and RollingDiffusion [32] are typical examples. Our ablation studies investigate the influence of adding noise to history frames.
Classifier-Free Guidance (CFG) over history frames applies different masks or noise levels to opposite sides of guidance to amplify the forgetting-drifting trade-off. HistoryGuidance [34] demonstrates this approach. Our ablation studies include guidance-based noise scheduling.
Anchor frames can be used as planning elements for video generation. StreamingT2V [15] and ART-V [43] use reference images as anchors. Video planning approaches [26, 59, 16, 48, 1, 49] use image or video anchors for content planning.
Compressing latent space can improve the efficiency of video diffusion models. LTXVideo [13] shows that a highly compressed latent space can be used for diffusing videos efficiently. PyramidFlow [20] diffuses video latents in a pyramid and renoise downsampled latents in that pyramid to reduce computation costs. FAR [12] proposes a multi-level causal attention structure to establish long-short-term causal context pacifying and KV caches. HiTVideo [62] uses hierarchical tokenizers to enhance the video generation with autoregressive language models.
The trade-off between forgetting and drifting is also evidenced by previous discussions. CausVid [53] shows that when the video generator is causal, the quality degradation appears at the end of the video and the final video length may be subject to an upper bound. DiffusionForcing [4] discussed that the cause of this drift may be related to error accumulation in models’ observation disparity between training and inference. Wang et al. [40] discussed that a model with stronger memory may suffer more from drifting and error accumulation.
2.2 Long Video Generation
Extending video generation beyond short clips remains an open problem. LVDM [14] generates long videos using latent diffusion, while Phenaki [38] creates variable-length videos from sequences of text prompts. Gen-L-Video [39] applies temporal co-denoising for multi-text conditioned videos, and FreeNoise [31] extends pretrained models without additional training via noise rescheduling. NUWA-XL [50] implements a Diffusion-over-Diffusion architecture with coarse-to-fine processing, while Video-Infinity [36] overcomes computational constraints through distributed generation. StreamingT2V [15] produces consistent, dynamic, and extendable videos without hard cuts, and CausVid [53] transforms bidirectional models into fast autoregressive ones through distillation. Recent advances include GPT-like architecture (ViD-GPT [11]), multi-event generation (MEVG [30]), attention control for multi-prompt generation (DiTCtrl [3]), precise temporal control (MinT [44]), history-based guidance (HistoryGuidance [34]), unified next-token and full-sequence diffusion (DiffusionForcing [4]), SpectralBlend temporal attention (FreeLong [27]), video autoregressive modeling (FAR [12]), and test-time training (TTT [6]).
2.3 Efficient Architectures for Video Generation
We discuss typical methods to improve video model efficiency. Linear attention [2, 47, 41, 5, 54, 21] reduces the attention complexity by reformulating linear operations. Sparse attention [45, 57, 58, 46] computes attention only on important token pairs, skipping negligible attention values. Low-bit computation [24, 60, 23] quantizes model weights and activations to lower precision for faster computation. Low-bit attention [56, 55] specifically optimizes attention computation through quantization techniques. Hidden state caching [29, 25] reuses intermediate computations across diffusion timesteps to avoid redundant calculations. Distillation [35, 28, 51, 52] transfers knowledge from larger models to smaller ones or reduces sampling steps while preserving quality.
3 Method
We consider a video generation model that predicts next frames repeatedly to form a video. For simplicity, we consider diffusion-based next-frame-section prediction models using Diffusion Transformers (DiTs) that generate a section of $S$ unknown frames $\bm{X}\in\mathbb{R}^{S\times h\times w\times c}$ conditioned on $T$ input frames $\bm{F}\in\mathbb{R}^{T\times h\times w\times c}$. All definitions of frames and pixels refer to latent representations, as most modern models operate in latent space.
For next-frame (or next-frame-section) prediction, $S$ is typically 1 (or a small number). We focus on the challenging case where $T\gg S$. With per-frame context length $L_{f}$ (typically $L_{f}\approx 1560$ for each 480p frame in Hunyuan/Wan/Flux), the vanilla DiT yields total context length $L=L_{f}(T+S)$. This causes context length explosion when $T$ is large.
3.1 FramePack
We observe that the input frames have different importance when predicting the next frame, and we can prioritize the input frames according to their importance. Without loss of generality, we consider a simple case where the temporal proximity reflects importance: frames temporally closer to the prediction target can be more relevant. We enumerate all frames with $\bm{F}_{0}$ being the most important (e.g., the most recent) and $\bm{F}_{T-1}$ being the least (e.g., the oldest).
We define a length function $\phi(\bm{F}_{i})$ that determines each frame’s context length after VAE encoding and transformer patchifying, applying progressive compression to less important frames as
$\phi(\bm{F}_{i})=\frac{L_{f}}{\lambda^{i}}\,,$ (1)
where $\lambda>1$ is a compression parameter. The frame-wise compression is achieved by manipulating the transformer’s patchify kernel size in the input layer (e.g., $\lambda=2,i=5$ means a kernel size where the product of all dims equals $322^{5}=$ 32 like the 3D kernel $2\times 4\times 4$, or $8\times 2\times 2$, etc.). The total context length then follows a geometric progression
$L=S\cdot L_{f}+L_{f}\cdot\sum_{i=0}^{T-1}\frac{1}{\lambda^{i}}=S\cdot L_{f}+L_{f}\cdot\frac{1-1/\lambda^{T}}{1-1/\lambda}\,,$ (2)
and when $T\to\infty$, the total context length converges to
$\lim_{T\to\infty}L=S\cdot L_{f}+\frac{1}{1-1/\lambda}\cdot L_{f}=\left(S+\frac{\lambda}{\lambda-1}\right)\cdot L_{f}\,,$ (3)
and this bounded context length makes FramePack’s computation bottleneck invariant to the input frame number $T$.
Since most hardware supports efficient matrix processing by powers of 2, we mainly discuss the case of $\lambda=2$ in this paper. Note that we can represent arbitrary compression rates by duplicating (or dropping) several specific terms in the power-of-2 sequence: considering the accumulation $\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=\frac{2}{2-1}=2$, if we want to loosen it a bit, for example to $2.625$, we can duplicate the terms $\frac{1}{2}$ and $\frac{1}{8}$ so that $\frac{1}{1}+\frac{1}{2}+(\frac{1}{2})+\frac{1}{4}+\frac{1}{8}+(\frac{1}{8})+\frac{1}{16}+...=\frac{1}{2}+\frac{1}{8}+\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=2.625$. Following this, one can cover arbitrary rates by converting the rate value to binary bits and then translate every bit.
The patchifying operations in most DiTs are 3D, and we denote the 3D kernel as $(p_{f},p_{h},p_{w})$ representing the steps in frame number, height, and width. A same compression rate can be achieved by multiple possible kernel sizes, e.g., the compression rate of $64$ can be achieved by $(1,8,8)$, $(4,4,4)$, $(16,2,2)$, $(64,1,1)$, etc. These would lead to different FramePack compression schedules.
Figure 1: Ablation Variants of FramePack. We present several typical kernel structures of FramePack with commonly used kernel sizes and compression rates. This list does not necessarily cover all popular variants, and more structures can be developed in a similar way.
Independent patchifying parameters
We observe that features from deep neural networks at different compression rates exhibit notable differences. Empirical evidence shows that using independent parameters for the different input projections at multiple compression rates facilitates stabilized learning. We assign the most commonly used input compression kernels as independent neural network layers: $(2,4,4)$, $(4,8,8)$, and $(8,16,16)$. For higher compressions (e.g., at $(16,32,32)$), we first downsample (e.g., with $2\times 2\times 2$) and then use the largest kernel $(8,16,16)$. This allows us to handle all compression rates. When training these new input projection layers, we initialize their weights by interpolating from the pretrained patchifying projection (e.g., the $(2,4,4)$ projection of HunyuanVideo/Wan).
Tail options
While in theory FramePack can process videos of arbitrary length with a fixed, invariant context length, practical considerations arise when the input frame length becomes extremely large. In the tail area, frames may fall below a minimum unit size (e.g., a single latent pixel). We discuss 3 options to handle the tail: (1) simply delete the tail; (2) allow each tail frame to increase the context length by a single latent pixel; (3) apply global average pooling to all tail frames and process them with the largest kernel. In our tests, the visual differences between these options are relatively negligible. We note that the tail refers to the least important frames, not always the oldest frames (in some cases, we can assign old frames with higher importance).
RoPE alignment
When encoding inputs with different compression kernels, the different context lengths require RoPE (Rotary Position Embedding) alignment. RoPE generates complex numbers with real and imaginary parts for each token position across all channels, which we refer to as “phase”. RoPE typically multiplies the phase to neural network features channel-wise. To match the RoPE encoding after compression, we directly downsample (using average pooling) the RoPE phases to match the compression kernels.
3.2 FramePack Variants
The above FramePack structure is discussed with the frame-wise importance defined by simple temporal proximity, and the compression schedule with simple geometric progression. We consider more variants for practical applications and optimal quality.
We visualize typical structures in Fig. 1. First, compression levels can be duplicated and combined with higher compression rates. In Fig. 1-(b), a power-of-4 sequence with each level repeated 3 times allows for same kernel sizes in frame width and height, making the compression more compact. Compression can also be applied in the temporal dimension (e.g., using a power-of-two sequence), as in Fig. 1-(c). This method encodes multiple frames in the same tensor, which is naturally aligned with DiT architectures.
We discuss alternative frame-wise importance modeling beyond simple temporal proximity in Fig. 1-(d,e). For instance, we can assign full-length context to the oldest frame, or treat both beginning and ending frames as equally important while applying higher compression to middle frames. These structures are particularly effective for applications like image-to-video generation, where user-provided initial frames carry higher importance.
3.3 Anti-drifting Sampling
Figure 2: Visualization of Sampling Approaches. We present sampling approaches to generate frames in different temporal orders. The shadowed squares are the generated frames in each iteration, whereas the white squares are the iteration inputs. Note that only the vanilla approach relies on causal inputs, and both the anti-drifting and inverted anti-drifting approaches receive bi-directional inputs.
Drifting is a common problem in next-frame prediction models where visual quality degrades as video length increases. While the underlying cause remains an open research problem, we observe that drifting only happens in causal sampling (i.e., when models only access past frames). We show that providing access to future frames (even a single future frame) will get rid of drifting. We point out that bi-directional context, rather than strictly causal dependencies, might be fundamental for maintaining video quality.
The vanilla sampling method shown in Fig. 2-(a), i.e., iteratively predicting future frames, can be modified into Fig. 2-(b), where the first iteration simultaneously generates both beginning and ending sections, while subsequent iterations fill the gaps between these anchors. This bi-directional approach prevents drifting since the ending frames are established in the first iteration, and all future generations attempt to approximate them.
We discuss an important variant by inverting the sampling order in Fig. 2-(b) into Fig. 2-(c). This approach is effective for image-to-video generation because it can treat the user input as a high-quality first frame, and continuously refines generations to approximate the user frame (which is unlike Fig. 2-(b) that does not approximate the first frame), leading to overall high-quality videos.
All three methods in Fig. 2-(a,b,c) can generate videos of arbitrary length. Method (a) achieves this through direct iterative generation, while in methods (b) and (c), we can dynamically move the ending sections (or generated frames) to greater distances as our generated frames approach them. Alternatively, in practice, setting a sufficiently large time range (e.g., 1 minute) in the first iteration typically satisfies practical requirements.
RoPE with random access
These sampling methods require modifications to RoPE to support non-consecutive phases (time indices of frames). This is achieved by skipping the non-queried phases (indices) in the time dimension.
4 Experiments
4.1 Ablative Naming
To simplify the presentation of the experiments, we use a common naming convention for all ablative structures. A FramePack name is represented as a string such as td_f16k4f4k2f1k1_g9_x_f1k1. We explain the meaning of this notation:
Kernel: A kernel name is like k1h2w2. The k stands for “kernel”, and k1h2w2 indicates a patchify kernel with shape $(1,2,2)$, where the temporal size is $1$, the height is $2$, and the width is $2$.
Kernel (simplified): For simplicity, since kernels that are multiples of $(1,2,2)$ are commonly used, we use abbreviated notation such as k1 that only denotes the temporal dimension. Specifically, k1 represents k1h2w2 (the kernel $(1,2,2)$), k2 represents k2h4w4 (the kernel $(2,4,4)$), k4 represents k4h8w8 (the kernel $(4,8,8)$), etc.
Encoding frames: The notation f16k4 indicates that $16$ frames are encoded by the kernel k4 (the simplified k4h8w8) with kernel size $(4,8,8)$.
Packing: The notation f16k4f4k2f1k1 shows a way to encode $19$ contiguous frames with three compression levels: the first $16$ frames are encoded by the kernel k4 (the kernel $(4,8,8)$), the next $4$ frames are encoded by the kernel k2 (the kernel $(2,4,4)$), and the last $1$ frame is encoded by the kernel k1 (the kernel $(1,2,2)$).
Tail: We append the notation with td, ta, or tc to indicate the tail frames before or after packing, such as td_f16k4f4k2f1k1. The three options are as discussed in Section 3.1. Herein, the “delete” option td deletes the tail. The “append” option ta compresses each tail frame by performing a 3D pooling of $(1,32,32)$ and then encodes with the nearest kernel, and the “compress” option tc uses global average pooling for all tail frames and compresses them with the nearest kernel.
Skipping: The notation x skips an arbitrary number of frames (including 0 frames).
Generating: The notation g9 means generating 9 frames.
With the above naming convention, we can represent all ablative structures in a compact form. Note that this naming also implies the sampling approach as discussed in Section 3.3. Consider the three similar structures and their sampling:
td_f16k4f4k2f1k1_g9: The vanilla sampling that generates frames in temporal order.
td_f16k4f4k2f1k1_g9_x_f1k1: The anti-drifting sampling with an endpoint frame.
f1k1_x_g9_f1k1f4k2f16k4_td: The inverted anti-drifting sampling in inverted temporal order.
4.2 Base Model and Implementation Details
We implement FramePack with Wan and HunyuanVideo. We implement both the text-to-video and image-to-video structures, though both are naturally supported by next-frame-section prediction models and do not need architecture modifications.
The Wan base model is the official Wan2.1 model. The HunyuanVideo model is an improved version of official HunyuanVideo models to match the capability of Wan2.1, with the modifications: (1) adding the SigLip-Vision model google/siglip-so400m-patch14-384 as a vision encoder, (2) removing the reliance on Tencent’s internal MLLM, (3) freezing LLama3.1 as a pure text model and ignoring its multi-modality, and (4) continued training on high-quality data.
Note that the numerical results presented in this paper should not be interpreted as direct comparisons between Wan and HunyuanVideo, as both were trained with similar computational budgets while Wan’s larger model size means fewer relative resources (e.g., batch size). Both models demonstrate comparable quality after sufficient training. We recommend HunyuanVideo as the default configuration for more efficient training and faster inference.
Dataset
We follow the guidelines of LTXVideo [13]’s dataset collection pipeline to gather data at multiple resolutions and quality levels. All data are filtered using quality measurements and motion scores to ensure a high-quality diverse distribution. We use aspect ratio bucketing for multi-resolution training with a minimum unit size of $32$ pixels. For example, the buckets at 480p resolution include: $(416,960)$, $(448,864)$, $(480,832)$, $(512,768)$, $(544,704)$, $(576,672)$, $(608,640)$, $(640,608)$, $(672,576)$, $(704,544)$, $(768,512)$, $(832,480)$, $(864,448)$, and $(960,416)$.
Small-scale Training
FramePack achieves a batch size of 64 on a single 8×A100-80G node with the 13B HunyuanVideo model at 480p resolution (without using any image-based workaround training that many community LoRAs are built on top of). This batch size is comparable even to image diffusion models like the 12B Flux, making FramePack suitable for personal or laboratory-scale training and experimentation.
Training
We conduct all experiments using H100 GPU clusters. Each FramePack variant in the ablation studies consumes approximately 48 hours of training, while the final models are trained for one week. We use the Adafactor optimizer with a fixed learning rate of 1e-5 without learning rate scheduling. The gradient norm clip is set to 0.5.
Lower flow shift
Since the next-frame-section prediction methods generate smaller 3D tensors at each inference compared to full-video diffusion, the models can be trained with much lower flow shift values. We use Flux’s dynamic flow shift to train all models. We find that this leads to sharp and clean results that are closer to real videos.
4.3 Evaluation Metrics
We discuss the metrics for evaluating ablative architectures. The tested inputs consist of 512 real user prompts for text-to-video and 512 image-prompt pairs for image-to-video tasks. All test samples were curated from real users to ensure diversity and real-world applicability. For quantitative tests, we by default use 30 seconds for long videos and 5 seconds for short videos.
4.3.1 Multi-dimension Metrics
Multiple metrics for video evaluations are consistent with common benchmarks, e.g., VBench [18], VBench2 [61], etc. All scores are normalized using the same numerical system with VBench.
Clarity: The MUSIQ [22] image quality predictor trained on the SPAQ [10] dataset. This metric measures artifacts such as noise and blurring in the generated frames.
Aesthetic: The LAION aesthetic predictor [33]. This metric measures the aesthetic values perceived with a CLIP-based estimator.
Motion: The video frame interpolation model [17] modified by VBench to measure the smoothness of motion.
Dynamic: The RAFT [37] modified by VBench to estimate the degree of dynamics. Note that the “dynamic” metric and “motion” metric represent a trade-off, e.g., a still image may rank high on motion smoothness but will be penalized by low dynamic degrees.
Semantic: The video-text score computed by ViCLIP [42]. This metric measures the overall semantic consistency between the generated video and the input text.
Anatomy: The ViT [9] pretrained by VBench for identifying the per-frame presence of hands, faces, bodies, etc.
Identity: The facial feature similarity using ArcFace [7] with face detection by RetinaFace [8] to measure the identity consistency.
Table 1: Ablation study. We evaluate different FramePack configurations across multiple global metrics, drifting metrics, and human assessments. The table is divided into three groups based on sampling approach: vanilla sampling (top), anti-drifting sampling (middle), and inverted anti-drifting sampling (bottom). The tests are conducted with HunyuanVideo as base. Bests in bold. ELO differences within ±$\pm 16$ are considered ties. Variant Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ td_f8k8f4k4f2k2f1k1_g9 67.33% 65.94% 94.53% 92.91% 20.06% 66.97% 71.72% 3.25% 3.45% 7.45% 16.56% 1084 8 td_f64k8f16k4f4k2f1k1_g9 67.71% 66.71% 95.09% 93.91% 21.39% 66.56% 71.50% 3.22% 3.38% 7.25% 16.43% 1068 9 td_f512k8f64k4f8k2f1k1_g9 67.74% 66.44% 94.27% 92.91% 21.90% 67.73% 71.35% 3.18% 3.32% 7.05% 15.95% 1088 8 td_f512k16f64k8f8k2f1k1_g9 66.13% 64.52% 94.77% 94.18% 19.21% 65.72% 71.51% 3.25% 3.45% 7.85% 17.62% 1072 9 td_f16k4f4k2f1k1_g9 67.37% 65.05% 95.97% 91.66% 19.15% 65.38% 71.73% 3.22% 3.48% 6.65% 17.36% 1076 9 td_f16k4f2k2f1k1_g1 65.81% 64.39% 94.91% 94.92% 19.20% 64.16% 69.70% 3.43% 3.77% 9.81% 20.89% 1036 11 td_f16k4f2k2f1k1_g4 66.57% 64.99% 94.00% 82.85% 19.40% 65.97% 69.06% 3.36% 3.68% 8.55% 19.09% 1052 10 td_f16k4f2k2f1k1_g9 67.15% 65.29% 94.08% 92.97% 20.73% 66.46% 71.08% 3.18% 3.42% 7.45% 18.05% 1080 9 tc_f16k4f2k2f1k1_g9 67.62% 65.07% 94.02% 90.33% 21.71% 71.70% 71.01% 3.15% 3.25% 7.25% 17.21% 1092 8 ta_f16k4f2k2f1k1_g9 67.02% 66.44% 94.11% 91.09% 21.48% 71.92% 72.18% 3.12% 3.18% 7.05% 17.66% 1096 8 td_f8k8f4k4f2k2f1k1_g9_x_f1k1 68.46% 66.95% 96.46% 75.55% 22.88% 85.10% 75.84% 2.95% 2.85% 5.95% 15.87% 1132 4 td_f64k8f16k4f4k2f1k1_g9_x_f1k1 68.21% 67.75% 95.74% 85.97% 22.79% 81.09% 76.29% 2.92% 2.98% 5.95% 15.99% 1136 4 td_f512k8f64k4f8k2f1k1_g9_x_f1k1 68.62% 67.72% 95.05% 76.50% 22.44% 82.01% 76.65% 2.88% 2.92% 5.75% 15.94% 1116 5 td_f512k16f64k8f8k2f1k1_g9_x_f1k1 68.35% 67.89% 97.73% 76.48% 22.75% 78.40% 76.16% 2.85% 2.85% 5.55% 14.04% 1120 5 td_f16k4f4k2f1k1_g9_x_f1k1 68.42% 67.59% 96.74% 74.37% 23.69% 79.14% 77.37% 2.82% 2.78% 5.35% 14.90% 1140 4 td_f16k4f2k2f1k1_g1_x_f1k1 65.32% 67.97% 95.26% 81.69% 19.97% 74.57% 77.93% 2.78% 2.72% 4.95% 14.80% 1084 7 td_f16k4f2k2f1k1_g4_x_f1k1 69.92% 67.49% 97.84% 71.90% 21.12% 74.84% 77.53% 2.75% 2.65% 4.95% 14.98% 1100 6 td_f16k4f2k2f1k1_g9_x_f1k1 69.51% 69.15% 96.97% 77.41% 23.03% 83.10% 69.25% 2.72% 2.58% 4.75% 13.73% 1144 4 tc_f16k4f2k2f1k1_g9_x_f1k1 69.62% 68.42% 96.45% 82.27% 23.08% 81.68% 69.08% 2.68% 2.52% 4.55% 13.54% 1148 4 ta_f16k4f2k2f1k1_g9_x_f1k1 69.21% 68.84% 97.87% 76.22% 22.77% 81.70% 75.23% 2.65% 2.45% 4.35% 13.76% 1152 4 f1k1_x_g9_f1k1f2k2f4k4f8k8_td 69.62% 67.87% 97.93% 88.79% 23.73% 86.99% 78.63% 2.55% 2.35% 4.15% 12.71% 1205 1 f1k1_x_g9_f1k1f4k2f16k4f64k8_td 69.56% 67.48% 97.42% 86.68% 24.48% 86.35% 78.06% 2.45% 2.25% 3.95% 12.52% 1209 1 f1k1_x_g9_f1k1f8k2f64k4f512k8_td 69.35% 67.89% 97.85% 88.21% 24.66% 76.99% 79.56% 2.35% 2.15% 3.75% 12.13% 1213 1 f1k1_x_g9_f1k1f8k2f64k8f512k16_td 69.20% 68.76% 99.11% 89.18% 24.35% 77.62% 79.88% 2.35% 2.05% 3.55% 11.10% 1239 1 f1k1_x_g9_f1k1f4k2f16k4_td 69.25% 67.40% 98.59% 89.01% 24.24% 77.31% 79.16% 2.45% 1.95% 3.35% 9.22% 1221 1 f1k1_x_g1_f1k1f2k2f16k4_td 69.74% 67.04% 98.09% 79.85% 24.89% 77.27% 80.86% 2.55% 2.15% 3.75% 11.37% 1148 3 f1k1_x_g4_f1k1f2k2f16k4_td 70.28% 68.11% 98.30% 79.45% 24.87% 77.95% 80.23% 2.45% 2.05% 3.75% 11.12% 1164 2 f1k1_x_g9_f1k1f2k2f16k4_td 70.73% 67.97% 98.11% 88.76% 25.79% 78.45% 84.01% 2.35% 1.95% 3.65% 11.98% 1225 1 f1k1_x_g9_f1k1f2k2f16k4_tc 70.48% 68.74% 98.16% 89.18% 27.01% 87.21% 81.04% 2.18% 1.85% 2.54% 11.71% 1229 1 f1k1_x_g9_f1k1f2k2f16k4_ta 70.81% 67.31% 98.15% 80.72% 25.60% 85.60% 82.76% 2.25% 1.77% 2.95% 9.85% 1233 1
4.3.2 Drifting Measurements
We discuss the methods to measure drifting. Several metrics from VBench-Long [19] already integrate long-range consistency measurements and may indicate drifting. We propose a more direct metric, namely start-end contrast, to measure drifting from multiple aspects.
We observe that when drifting occurs, a significant difference emerges between the beginning and ending portions of a video across various quality metrics. We define the start-end contrast $\Delta_{\text{drift}}^{M}$ for an arbitrary quality metric $M$ as:
$\Delta_{\text{drift}}^{M}(V)=\left|M(V_{\text{start}})-M(V_{\text{end}})\right|\,,$ (4)
where $V$ is the tested video, $V_{\text{start}}$ represents the first 15% of frames, and $V_{\text{end}}$ represents the last 15% of frames. This start-end contrast can be applied to different metrics $M$ (e.g., motion score, image quality, etc.). The magnitude of $\Delta_{\text{drift}}^{M}(V)$ directly indicates the severity of drifting. Since video models may generate frames in different temporal orders (either forward or backward), we use the absolute difference to ensure our metric remains direction-agnostic.
4.3.3 Human Assessments
We collect human preferences from A/B tests. Each ablative architecture yields 100 results. The A/B tests are randomly distributed among ablations, and we ensure that each ablation covers at least 100 assessments. We report ELO-K32 score and the relative ranking.
4.4 Ablative Results
As shown in Table 1, we note several discoveries: (1) The inverted anti-drifting sampling method achieves the best results in 5 out of 7 metrics, while other sampling methods achieve at most a single best metric. (2) The inverted anti-drifting sampling achieves the best performance in all drifting metrics. (3) Human evaluations indicate that generating 9 frames per section yields better perception than generating 1 or 4 frames, as evidenced by the higher ELO scores for configurations with g9 compared to their g1 or g4 counterparts. (4) While vanilla sampling achieved the highest dynamic score, this is likely attributable to drifting effects rather than genuine quality. (5) We also observe that differences between specific configuration options within the same sampling approach are relatively small and random, suggesting that the sampling category contributes more to the overall performance difference.
4.5 Comparison to Alternative Architectures
We discuss several relevant alternatives to generate videos in various ways. The involved methods either enable longer video generation, reduce computational bottlenecks, or both. To be specific, we implement these variants on top of HunyuanVideo default architecture (33 latent frames) using a simple naive sliding window with half context length for history inputs.
Repeating image-to-video: Directly repeat the image-to-video inference to make longer videos.
Anchor frames: Use an image as the anchor frame to avoid drifting. We implement a structure that resembles StreamingT2V [15].
Causal attention: Finetune full attention into causal attention for easier KV cache and faster inference. We implement a structure that resembles CausVid [53].
Noisy history: Delay the denoising timestep on history latents so that history latents are noisy (but less noisy than the current generating latents). Reducing the reliance on the history is beneficial for interrupting error accumulation, thus mitigates drifting, but at the cost of aggravating forgetting. We implement a structure that resembles DiffusionForcing [4].
History guidance: Delay the denoising timestep on history latents but also put the completely noised history on the unconditional side of CFG guidance. This will speed up error accumulation thus aggravating drifting, but also enhance memory to mitigate forgetting. We implement a structure that resembles HistoryGuidance [34].
Table 2: Comparison with relevant architectures. We compare with other relevant methods across the same global metrics, drifting metrics, and human assessments used in the ablation study. The tests are conducted with HunyuanVideo as base. Bests in bold. ELO differences within ±$\pm 16$ are considered ties. Method Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ Repeating image-to-video 56.73% 56.15% 94.34% 65.56% 17.74% 69.41% 73.06% 9.51% 3.92% 9.95% 19.88% 1016 5 Anchor frames (resembling StreamingT2V [15]) 69.58% 67.35% 99.96% 74.97% 25.76% 85.01% 79.52% 2.85% 2.15% 3.45% 9.25% 1172 2 Causal attention (resembling CausVid [53]) 62.88% 59.41% 96.98% 67.56% 19.15% 72.74% 75.32% 7.45% 3.15% 6.75% 15.96% 1087 4 Noisy history (resembling DiffusionForcing [4]) 66.75% 63.78% 97.20% 91.45% 22.98% 77.60% 77.92% 4.35% 2.05% 3.15% 9.85% 1146 3 History guidance (resembling HistoryGuidance [34]) 68.05% 68.74% 97.01% 73.39% 24.88% 81.84% 83.42% 7.35% 2.21% 5.25% 12.78% 1151 3 Proposed (f1k1_x_g9_f1k1f2k2f16k4) 71.15% 68.71% 99.45% 89.29% 28.15% 86.53% 82.11% 2.25% 1.85% 2.68% 8.58% 1221 1
As shown in Table 2, we observe several findings: (1) The proposed method achieves the best results in 3 global metrics, while other methods excel in at most one or two metrics. (2) The proposed method achieves the best results across all drifting metrics. (3) This evaluation aligns with human perceptions as evidenced by the ELO score.
5 Conclusion
In this paper, we presented FramePack, a neural network structure that aims to address the forgetting-drifting dilemma in next-frame prediction models for video generation. FramePack applies progressive compression to input frames based on their importance, ensuring the total context length converges to a fixed upper bound regardless of video duration. This is achieved through manipulating transformer patchify kernel sizes for different compression rates. Combined with the anti-drifting sampling methods that incorporate bi-directional context through early-established endpoints or inverted temporal ordering, our approach enables longer video generation while maintaining unchanged computational bottlenecks. Experiments suggest that FramePack can improve model responsiveness and visual quality in inference, while allowing for higher batch sizes in training. The approach is compatible with existing video diffusion models and supports various compression variants that can be optimized for wider applications.

### v3

@@
-Packing Input Frame Context in Next-Frame Prediction Models for Video Generation
+Frame Context Packing and Drift Prevention in Next-Frame-Prediction Video Diffusion Models
 Abstract
 We present a neural network structure, FramePack, to train next-frame (or next-frame-section) prediction models for video generation.
-The FramePack compresses input frames to make the transformer context length a fixed number regardless of the video length.
-As a result, we are able to process a large number of frames using video diffusion with computation bottleneck similar to image diffusion.
-This also makes the training video batch sizes significantly higher (batch sizes become comparable to image diffusion training).
-We also propose an anti-drifting sampling method that generates frames in inverted temporal order with early-established endpoints to avoid exposure bias (error accumulation over iterations).
-Finally, we show that existing video diffusion models can be finetuned with FramePack, and their visual quality may be improved because the next-frame prediction supports more balanced diffusion schedulers with less extreme flow shift timesteps.
+FramePack compresses input frame contexts with frame-wise importance so that more frames can be encoded within a fixed context length, with more important frames having longer contexts.
+The frame importance can be measured using time proximity, feature similarity, or hybrid metrics.
+The packing method allows for inference with thousands of frames and training with relatively large batch sizes.
+We also present drift prevention methods to address observation bias (error accumulation), including early-established endpoints, adjusted sampling orders, and discrete history representation.
+Ablation studies validate the effectiveness of the anti-drifting methods in both single-directional video streaming and bi-directional video generation.
+Finally, we show that existing video diffusion models can be finetuned with FramePack, and analyze the differences between different packing schedules.
 1 Introduction
-Forgetting and drifting are the two most critical problems in next-frame (or next-frame-section) prediction models for video generation.
-Herein, "forgetting" refers to the fading of memory as the model struggles to remember earlier content and maintain consistent temporal dependencies, whereas "drifting" refers to the iterative degradation of visual quality due to error accumulation over time (also called exposure bias).
-A fundamental dilemma emerges when attempting to simultaneously address both forgetting and drifting: any method that mitigates forgetting by enhancing memory may also make error accumulation/propagation faster, thereby exacerbating drifting; any method that reduces drifting by interrupting error propagation and weakening the temporal dependencies (e.g., masking or re-noising the history) may also worsen the forgetting.
-This essential trade-off hinders the scalability of next-frame prediction models.
-The forgetting problem leads to a naive solution to encode more frames, but this quickly becomes computationally intractable due to the quadratic attention complexity of transformers (or sub-quadratic optimizations like FlashAttn, etc.).
-Moreover, video frames contain significant temporal redundancy, rendering naive full-context approaches less efficient.
-The substantial duplication of visual features across consecutive frames reveals the potential to design effective compression systems to facilitate memorization.
-The drifting problem is influenced by memorizing mechanisms from multiple aspects.
-The source of drifting lies in initial errors that occur in individual frames, while the effect is the propagation and accumulation of these errors across subsequent frames, leading to degraded visual quality.
-On one hand, a stronger memorizing mechanism leads to better temporal consistency and reduces the occurrence of initial errors, thereby mitigating drifting.
-On the other hand, a stronger memorizing mechanism also memorizes more errors and thus accelerates error propagation when errors do occur, exacerbating drifting.
+Forgetting and drifting are the two most critical problems in next-frame-prediction models for video generation.
+“Forgetting” refers to the fading of memory as the model struggles to remember earlier content and maintain consistent temporal dependencies.
+“Drifting” refers to the degradation of visual quality due to error accumulation over time (also called exposure/observation bias).
+A fundamental dilemma emerges when attempting to simultaneously address both forgetting and drifting: any method that mitigates forgetting by enhancing memory may also increase error accumulation/propagation, thereby exacerbating drifting; any method that reduces drifting by interrupting error propagation needs to weaken temporal dependencies (e.g., masking or re-noising the history), thus worsening the forgetting.
+This is a fundamental trade-off that hinders the scalability of next-frame prediction models.
+A naive solution to forgetting is to encode more frames.
+But this approach quickly becomes computationally intractable due to the quadratic attention complexity of transformers (even with optimizations like Flash Attention [10], etc.).
+Moreover, video frames contain significant temporal redundancy, making naive full-context approaches very inefficient.
+The substantial duplication of visual features across consecutive frames suggests the potential to design effective compression systems to facilitate memorization.
+Drifting is influenced by memorizing mechanisms in several ways.
+The source of drifting lies in the initial errors that occur in individual frames, while the effect is the propagation and accumulation of these errors across subsequent frames, and the eventual drifting out of the train distribution.
+A stronger memorization mechanism, on one hand, can lead to better temporal consistency and reduce the occurrence of initial errors.
+On the other hand, it also memorizes more errors and thus accelerates error propagation when errors do occur.
 This paradoxical relationship between memory mechanisms and drifting necessitates carefully designed training/sampling methods to facilitate error correction or interrupt error propagation.
-In this paper, we propose FramePack as an anti-forgetting memory structure along with anti-drifting sampling methods.
-The FramePack structure addresses the forgetting problem by compressing input frames based on their relative importance, ensuring the total transformer context length converges to a fixed upper bound regardless of video duration.
-This enables the model to encode significantly more frames without increasing the computational bottleneck, thereby facilitating anti-forgetting.
-Moreover, we propose anti-drifting sampling methods that break the causal prediction chain and incorporate bi-directional context.
-These methods include generating endpoint frames before filling intermediate content, and an inverted temporal sampling approach where frames are generated in reverse order, each attempting to approach a known high-quality frame (e.g., an input frame in image-to-video tasks).
+In this paper, we propose FramePack as an anti-forgetting memory structure along with anti-drifting sampling and training methods.
+The FramePack structure addresses the forgetting problem by compressing input frames based on their relative importance, ensuring that the total transformer context length converges to a fixed upper bound.
+We consider both time-proximity-based and feature-similarity-based importance measures.
+Afterwards, we propose anti-drifting sampling methods that break the causal prediction chain and incorporate bi-directional contexts by planning single or multiple endpoint frames.
+We also present an anti-drifting training method to convert frame history into discrete tokens so as to reduce the history disparity between training and inference.
 We show that these methods effectively reduce the occurrence of errors and prevent their propagation.
-We demonstrate that existing pretrained video diffusion models (e.g., HunyuanVideo, Wan, etc.) can be finetuned with FramePack.
+We demonstrate that existing pretrained video diffusion models (e.g., HunyuanVideo [28], Wan [48], etc.) can be finetuned with FramePack.
 Our experiments reveal several findings: because next-frame prediction generates smaller tensor sizes per step compared to full-video generation, it enables more balanced diffusion schedulers with less extreme flow shift timesteps.
-We observe that the resulting less aggressive schedulers may lead to improved visual quality, beyond the direct goals of addressing the forgetting and drifting problems.
+We also show that the efficient implementations of FramePack can process thousands of frames with 13B models even on laptops (e.g., 6GB or 8GB GPU memory).
 2 Related Work
 2.1 Anti-forgetting and Anti-drifting
-Noise scheduling and augmentation in history frames modify noise levels at specific timesteps, video times, or image frequencies to create causal computation or anti-drifting effects.
-These methods generally reduce dependency on past frames.
+The trade-off between forgetting and drifting is also evidenced by previous discussions.
+CausVid [65] shows that when the video generator is causal, the quality degradation appears at the end of the video and the high-quality part may be subject to an upper bound length.
+DiffusionForcing [6] discussed that the cause of this drift may be related to error accumulation in models’ observation disparity between training and inference.
+Wang et al[51] discussed that a model with stronger memory may suffer more from drifting and error accumulation.
+Noise scheduling and augmentation in history frames modify noise levels at specific timesteps, video times, or image frequencies to mitigate drifting.
+These methods generally reduce the dependency on past frames.
 DiffusionForcing [6] and RollingDiffusion [39] are typical examples.
@@
 Compressing latent space can improve the efficiency of video diffusion models.
+FlexTok [2] adjusts token context length to achieve different levels of visual content compression.
 LTXVideo [17] shows that a highly compressed latent space can be used for diffusing videos efficiently.
-PyramidFlow [20] diffuses video latents in a pyramid and renoise downsampled latents in that pyramid to reduce computation costs.
+PyramidFlow [25] diffuses video latents in a pyramid and re-noises downsampled latents in that pyramid to reduce computation costs.
 FAR [16] proposes a multi-level causal attention structure to establish long-short-term causal context pacifying and KV caches.
 HiTVideo [76] uses hierarchical tokenizers to enhance the video generation with autoregressive language models.
-The trade-off between forgetting and drifting is also evidenced by previous discussions.
-CausVid [53] shows that when the video generator is causal, the quality degradation appears at the end of the video and the final video length may be subject to an upper bound.
-DiffusionForcing [4] discussed that the cause of this drift may be related to error accumulation in models’ observation disparity between training and inference.
-Wang et al.
-[40] discussed that a model with stronger memory may suffer more from drifting and error accumulation.
+Memory in world models often involves different modeling of long-term memory.
+Typical examples are 3D geometry like mesh and proxy [38, 50, 67, 66].
+Training diffusion models with domain data can also bake the memory into the model, with full model training [1, 46] or low-rank methods [21].
+Retrieval-based memory mechanisms like WorldMem [58] are efficient when the task prioritizes reconstructing history contents.
 2.2 Long Video Generation
@@
 NUWA-XL [62] implements a Diffusion-over-Diffusion architecture with coarse-to-fine processing, while Video-Infinity [44] overcomes computational constraints through distributed generation.
-StreamingT2V [15] produces consistent, dynamic, and extendable videos without hard cuts, and CausVid [53] transforms bidirectional models into fast autoregressive ones through distillation.
+StreamingT2V [20] produces consistent, dynamic, and extendable videos without hard cuts, and CausVid [65] transforms bidirectional models into fast autoregressive models through distillation.
 Recent advances include GPT-like architecture (ViD-GPT [15]), multi-event generation (MEVG [36]), attention control for multi-prompt generation (DiTCtrl [5]), precise temporal control (MinT [55]), history-based guidance (HistoryGuidance [41]), unified next-token and full-sequence diffusion (DiffusionForcing [6]), SpectralBlend temporal attention (FreeLong [33]), video autoregressive modeling (FAR [16]), and test-time training (TTT [9]).
-2.3 Efficient Architectures for Video Generation
-We discuss typical methods to improve video model efficiency.
-Linear attention [2, 47, 41, 5, 54, 21] reduces the attention complexity by reformulating linear operations.
-Sparse attention [45, 57, 58, 46] computes attention only on important token pairs, skipping negligible attention values.
-Low-bit computation [24, 60, 23] quantizes model weights and activations to lower precision for faster computation.
-Low-bit attention [56, 55] specifically optimizes attention computation through quantization techniques.
-Hidden state caching [29, 25] reuses intermediate computations across diffusion timesteps to avoid redundant calculations.
-Distillation [35, 28, 51, 52] transfers knowledge from larger models to smaller ones or reduces sampling steps while preserving quality.
-3 Method
+Harvey et al[18] proposes a flexible approach for modeling long contexts with dilatation (Hierarchy-2) and propagation.
+Generating longer videos often requires efficient architectures, e.g., linear attention [4, 59, 52, 8, 68, 26], sparse attention [56, 71, 72, 57], low-bit computation [30, 74, 29], low-bit attention [70, 69], hidden state caching [35, 31], distillation [42, 34, 63, 64], etc.
+3 Packing Frame Context
 We consider a video generation model that predicts next frames repeatedly to form a video.
-For simplicity, we consider diffusion-based next-frame-section prediction models using Diffusion Transformers (DiTs) that generate a section of $S$ unknown frames $\bm{X}\in\mathbb{R}^{S\times h\times w\times c}$ conditioned on $T$ input frames $\bm{F}\in\mathbb{R}^{T\times h\times w\times c}$.
+For simplicity, we consider next-frame-section prediction models using Diffusion Transformers (DiTs) that generate a section $\bm{X}$ of $S$ unknown frames so that $\bm{X}\in\mathbb{R}^{S\times h\times w\times c}$, conditioned on a section $\bm{F}$ of $T$ input frames so that $\bm{F}\in\mathbb{R}^{T\times h\times w\times c}$.
 All definitions of frames and pixels refer to latent representations, as most modern models operate in latent space.
@@
 With per-frame context length $L_{f}$ (typically $L_{f}\approx 1560$ for each 480p frame in Hunyuan/Wan/Flux), the vanilla DiT yields total context length $L=L_{f}(T+S)$.
-This causes context length explosion when $T$ is large.
-3.1 FramePack
-We observe that the input frames have different importance when predicting the next frame, and we can prioritize the input frames according to their importance.
-Without loss of generality, we consider a simple case where the temporal proximity reflects importance: frames temporally closer to the prediction target can be more relevant.
-We enumerate all frames with $\bm{F}_{0}$ being the most important (e.g., the most recent) and $\bm{F}_{T-1}$ being the least (e.g., the oldest).
-We define a length function $\phi(\bm{F}_{i})$ that determines each frame’s context length after VAE encoding and transformer patchifying, applying progressive compression to less important frames as
+This causes a context length explosion when $T$ is large.
+We observe that the input frames have different importance when predicting the next frame, and we can prioritize them according to their importance.
+3.1 Time Proximity Based Packing
+We first consider a baseline case where the temporal proximity reflects frame importance (Fig.
+1-(a)).
+More advanced cases involving similarity-based importance are covered in §3.2.
+With frames temporally closer to the prediction target being more relevant, we enumerate all frames with $\bm{F}_{0}$ being the most important (e.g., the most recent) and $\bm{F}_{T-1}$ being the least (e.g., the oldest).
+We define a length function $\phi(\bm{F}_{i})$ that determines each frame’s context length after VAE encoding and transformer patchifying by applying progressive compression
 $\phi(\bm{F}_{i})=\frac{L_{f}}{\lambda^{i}}\,$, (1)
 where $\lambda>1$ is a compression parameter.
-The frame-wise compression is achieved by manipulating the transformer’s patchify kernel size in the input layer (e.g., $\lambda=2,i=5$ means a kernel size where the product of all dims equals $322^{5}=$ 32 like the 3D kernel $2\times 4\times 4$, or $8\times 2\times 2$, etc.).
+The frame-wise compression is achieved by manipulating the transformer’s patchify kernel size in the input layer (e.g., $\lambda=2,i=5$ means a kernel size where the product of all dims equals $2^{5}=32$ like the 3D kernel $2\times 4\times 4$, or $8\times 2\times 2$, etc.).
 The total context length then follows a geometric progression
 $L=S\cdot L_{f}+L_{f}\cdot\sum_{i=0}^{T-1}\frac{1}{\lambda^{i}}=S\cdot L_{f}+L_{f}\cdot\frac{1-1/\lambda^{T}}{1-1/\lambda}\,$, (2)
-and when $T\to\infty$, the total context length converges to
-$\lim_{T\to\infty}L=S\cdot L_{f}+\frac{1}{1-1/\lambda}\cdot L_{f}=\left(S+\frac{\lambda}{\lambda-1}\right)\cdot L_{f}\,,$ (3)
-and this bounded context length makes FramePack’s computation bottleneck invariant to the input frame number $T$.
+and when $T\to\infty$, the total context length converges to $\lim_{T\to\infty}L=(S+\frac{\lambda}{\lambda-1})$.
+This bounded context length makes FramePack’s compression bottleneck invariant to the input frame number $T$.
 Since most hardware supports efficient matrix processing by powers of 2, we mainly discuss the case of $\lambda=2$ in this paper.
 Note that we can represent arbitrary compression rates by duplicating (or dropping) several specific terms in the power-of-2 sequence: considering the accumulation $\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=\frac{2}{2-1}=2$, if we want to loosen it a bit, for example to $2.625$, we can duplicate the terms $\frac{1}{2}$ and $\frac{1}{8}$ so that $\frac{1}{1}+\frac{1}{2}+(\frac{1}{2})+\frac{1}{4}+\frac{1}{8}+(\frac{1}{8})+\frac{1}{16}+...=\frac{1}{2}+\frac{1}{8}+\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=2.625$.
-Following this, one can cover arbitrary rates by converting the rate value to binary bits and then translate every bit.
+Following this, one can cover arbitrary rates by converting the rate value to binary bits and then translating every bit.
+Packing schedules
+Figure 1: FramePack variants.
+We present frame packing methods using time proximity or feature similarity.
+We discuss several typical kernel structures.
+This list does not necessarily cover all popular variants, and more structures can be developed in a similar way.
 The patchifying operations in most DiTs are 3D, and we denote the 3D kernel as $(p_{f},p_{h},p_{w})$ representing the steps in frame number, height, and width.
 A same compression rate can be achieved by multiple possible kernel sizes, e.g., the compression rate of $64$ can be achieved by $(1,8,8)$, $(4,4,4)$, $(16,2,2)$, $(64,1,1)$, etc.
-These would lead to different FramePack compression schedules.
-Figure 1: Ablation Variants of FramePack.
-We present several typical kernel structures of FramePack with commonly used kernel sizes and compression rates.
-This list does not necessarily cover all popular variants, and more structures can be developed in a similar way.
+Compression levels can be duplicated and combined with higher compression rates.
+We discuss more packing ways in Fig.
+1-(c): duplication allows for same kernel sizes in frame width and height, making the compression more compact; temporal kernel can compress contiguous frames into a single tensor; symmetric progression treats both beginning and ending frames as equally important.
 Independent patchifying parameters
-We observe that features from deep neural networks at different compression rates exhibit notable differences.
 Empirical evidence shows that using independent parameters for the different input projections at multiple compression rates facilitates stabilized learning.
@@
 For higher compressions (e.g., at $(16,32,32)$), we first downsample (e.g., with $2\times 2\times 2$) and then use the largest kernel $(8,16,16)$.
-This allows us to handle all compression rates.
-When training these new input projection layers, we initialize their weights by interpolating from the pretrained patchifying projection (e.g., the $(2,4,4)$ projection of HunyuanVideo/Wan).
+We initialize their separated weights by interpolating from the pretrained patchifying projection (e.g., the $(2,4,4)$ projection of HunyuanVideo/Wan).
 Tail options
-While in theory FramePack can process videos of arbitrary length with a fixed, invariant context length, practical considerations arise when the input frame length becomes extremely large.
-In the tail area, frames may fall below a minimum unit size (e.g., a single latent pixel).
-We discuss 3 options to handle the tail: (1) simply delete the tail; (2) allow each tail frame to increase the context length by a single latent pixel; (3) apply global average pooling to all tail frames and process them with the largest kernel.
+While in theory FramePack can process videos of arbitrary length with a fixed, invariant context length, frames may fall below a minimum unit size (e.g., a single latent pixel) when the input length becomes extremely large.
+We discuss 3 options to process the tail frames: (1) simply delete the tail; (2) allow each tail frame to increase the context length by a single latent pixel; (3) apply global average pooling to all tail frames and process them with the last kernel.
 In our tests, the visual differences between these options are relatively negligible.
-We note that the tail refers to the least important frames, not always the oldest frames (in some cases, we can assign old frames with higher importance).
 RoPE alignment
-When encoding inputs with different compression kernels, the different context lengths require RoPE (Rotary Position Embedding) alignment.
+When encoding inputs with different compression kernels, the different context lengths require RoPE (Rotary Position Embedding) [43] alignment.
 RoPE generates complex numbers with real and imaginary parts for each token position across all channels, which we refer to as “phase”.
-RoPE typically multiplies the phase to neural network features channel-wise.
-To match the RoPE encoding after compression, we directly downsample (using average pooling) the RoPE phases to match the compression kernels.
-3.2 FramePack Variants
-The above FramePack structure is discussed with the frame-wise importance defined by simple temporal proximity, and the compression schedule with simple geometric progression.
-We consider more variants for practical applications and optimal quality.
-We visualize typical structures in Fig.
-1.
-First, compression levels can be duplicated and combined with higher compression rates.
-In Fig.
-1-(b), a power-of-4 sequence with each level repeated 3 times allows for same kernel sizes in frame width and height, making the compression more compact.
-Compression can also be applied in the temporal dimension (e.g., using a power-of-two sequence), as in Fig.
-1-(c).
-This method encodes multiple frames in the same tensor, which is naturally aligned with DiT architectures.
-We discuss alternative frame-wise importance modeling beyond simple temporal proximity in Fig.
-1-(d,e).
-For instance, we can assign full-length context to the oldest frame, or treat both beginning and ending frames as equally important while applying higher compression to middle frames.
-These structures are particularly effective for applications like image-to-video generation, where user-provided initial frames carry higher importance.
-3.3 Anti-drifting Sampling
-Figure 2: Visualization of Sampling Approaches.
+We directly downsample (using average pooling) such phases to match the compression kernels.
+3.2 Feature Similarity Based Packing and Hybrid Approach
+The aforementioned frame ordering $\bm{F}_{0...T-1}$ can be seen as a result of sorting all history frames using their time positions.
+We note that such sorting can also use other metrics like feature similarity (Fig.
+1-(b)).
+For instance, consider a typical cosine similarity
+$\operatorname{sim}_{\cos}(\bm{F}_{i},\hat{\bm{X}})=\sum_{p}\frac{(\bm{F}_{i})_{p}\cdot\hat{\bm{X}}_{p}^{\top}}{\|(\bm{F}_{i})_{p}\|\|\hat{\bm{X}}_{p}\|}\,$, (3)
+where the sum is taken over pixels $p$.
+This measures the cosine similarity between each history frame and the estimated next frame section.
+Sorting the history frames using $\operatorname{sim}_{\cos}(\cdot)$ will produce a permutation $\bm{F}_{0...T-1}$ with $\bm{F}_{0}$ being the most similar and $\bm{F}_{T-1}$ being the least.
+This permutation can directly replace the aforementioned time proximity.
+Since similarity-based permutation may change abruptly when processing contiguous frames, we also consider a smooth time proximity modeling
+$\operatorname{sim}_{time}(\bm{F}_{i},\hat{\bm{X}})=e^{-(\operatorname{time}(\bm{F}_{i})-\operatorname{time}(\hat{\bm{X}}))^{2}}\,$, (4)
+where $\operatorname{time}(\cdot)$ gets the frames’ starting time measured in seconds. Consider the weighting
+[…]
+where $\lambda_{time}$ is a weighting parameter.
+Sorting the history frames using $\operatorname{sim}_{\text{hybrid}}(\cdot)$ will produce a permutation that transits relatively smoothly as the generating window moves forward in time.
+This hybrid approach is suitable for world model datasets (mainly video games) that require returning to previously visited views of scenes or events.
+Similar sorting can also be applied to facial identity metrics to facilitate movie generation applications that emphasize consistent human actors.
+This method will be evaluated in ablation experiments in the supplementary materials.
+4 Drift Prevention
+Drifting is a common problem in next-frame prediction models where visual quality degrades as video length increases.
+We discuss anti-drifting approaches by adjusting the sampling processes and history representations as shown in Fig.
+2.
+Figure 2: Anti-drifting sampling and training methods.
 We present sampling approaches to generate frames in different temporal orders.
 The shadowed squares are the generated frames in each iteration, whereas the white squares are the iteration inputs.
-Note that only the vanilla approach relies on causal inputs, and both the anti-drifting and inverted anti-drifting approaches receive bi-directional inputs.
-Drifting is a common problem in next-frame prediction models where visual quality degrades as video length increases.
-While the underlying cause remains an open research problem, we observe that drifting only happens in causal sampling (i.e., when models only access past frames).
-We show that providing access to future frames (even a single future frame) will get rid of drifting.
-We point out that bi-directional context, rather than strictly causal dependencies, might be fundamental for maintaining video quality.
+We also discuss the method to convert the frame history into a discrete representation.
+4.1 Planned Endpoints and Adjusted Sampling Order
+One possible explanation of drifting is that the modeling of the chained conditional probability $\mathbb{P}(\bm{X}_{t}|\bm{X}_{t-1})$ fails to approximate $\mathbb{P}(\bm{X}_{t})$ due to imperfect estimations.
+This perspective indicates that a strict causal system is more vulnerable to drifting than bi-directional models that directly approximate $\mathbb{P}(\bm{X}_{t}|\bm{X}_{t_{1}},\bm{X}_{t_{2}})$ where $t_{1}<t<t_{2}$.
+Endpoint planning
 The vanilla sampling method shown in Fig.
-2-(a), i.e., iteratively predicting future frames, can be modified into Fig.
+2-(a) can be modified into Fig.
 2-(b), where the first iteration simultaneously generates both beginning and ending sections, while subsequent iterations fill the gaps between these anchors.
-This bi-directional approach prevents drifting since the ending frames are established in the first iteration, and all future generations attempt to approximate them.
-We discuss an important variant by inverting the sampling order in Fig.
+This simple method can get rid of drifting in specific cases when the video motions are in a relatively small range, or when the motion patterns are repeated or periodic (e.g., dancing, talking, spinning, etc.), or when the motion content follows some texture patterns (e.g., fire flame, water flow, etc.).
+Inverted sampling (image-to-video)
+A variant by inverting the sampling order in Fig.
 2-(b) into Fig.
-2-(c).
-This approach is effective for image-to-video generation because it can treat the user input as a high-quality first frame, and continuously refines generations to approximate the user frame (which is unlike Fig.
-2-(b) that does not approximate the first frame), leading to overall high-quality videos.
-All three methods in Fig.
-2-(a,b,c) can generate videos of arbitrary length.
-Method (a) achieves this through direct iterative generation, while in methods (b) and (c), we can dynamically move the ending sections (or generated frames) to greater distances as our generated frames approach them.
-Alternatively, in practice, setting a sufficiently large time range (e.g., 1 minute) in the first iteration typically satisfies practical requirements.
+2-(c) is effective for image-to-video generation.
+In image-to-video, the first frame is a ground-truth user input, whereas the last frame is a generated endpoint that is not guaranteed to perfectly preserve the quality of the user input.
+All generations in Fig.
+2-(c) keep the direction to approximate the high-quality user input, leading to iteratively refined generations.
+Multiple endpoints
+The endpoint planning can be repeated with different prompts before filling in the gaps, resulting in a planned sequence of generation (Fig.
+2-(d)).
+Note that though the drifting might still happen in the endpoint-wise, in certain cases, when the prompted sections are distant enough, the error accumulation becomes almost negligible.
+This method is more flexible than single endpoint planning and supports more dynamic motions and more complicated storytelling.
 RoPE with random access
-These sampling methods require modifications to RoPE to support non-consecutive phases (time indices of frames). This is achieved by skipping the non-queried phases (indices) in the time dimension.
+These sampling methods require modifications to RoPE to support non-consecutive phases (time indices of frames). This is achieved by skipping the blank phases (indices) in the time dimension.
+4.2 History Discretization
+Another potential cause of drifting is the difference between training and inference distributions over the history frames.
+This indicates that drifting can be mitigated if the history representation is not sensitive enough to distinguish between the training and inference frames.
+Discrete integer tokens are well-suited to reduce the mode gap between training and inference distributions.
+This perspective is supported by empirical evidence that discrete autoregressive systems (e.g., LLMs) often demonstrate less obvious drifting than continuous autoregressive systems for visual content diffusion.
+We discretize the history as in Fig.
+2-(e).
+Consider a dataset $\Phi\in\mathbb{R}^{(B\times T\times H\times W)\times C}$ of precomputed latent videos, a $K$-Mean over all latent pixels will yield a codebook $\Omega\in\mathbb{R}^{K\times C}$ where $K\in\mathbb{Z}^{+}$ is an adjustable number.
+Any latent frame $\bm{F}\in\mathbb{R}^{T\times H\times W\times C}$ can be quantized by $Q(\cdot)$ with
+[…]
+where $p$ is pixel position, and $Q(\bm{F})\in[0,K-1]^{T\times H\times W}$ is a matrix of indices over the codebook $\Omega$.
+The matrix of indices can be converted back to latent videos by $\Omega_{Q(\bm{F})}\in\mathbb{R}^{T\times H\times W\times C}$.
+We replace all history frames $\bm{F}$ with $\Omega_{Q(\bm{F})}$ during training.
+Intuitively, when $K=1$, the history becomes meaningless as one single color, and the drifting is eliminated at the cost of giving up memory (errors do not accumulate but sections become unrelated); when $K\to\infty$, the effect is equivalent to no discretization, and the drifting remains.
+We show with ablation study that a suitable $K$ can minimize the error propagation while simultaneously producing plausible consistency between sections.
 5 Experiments
@@
 Kernel (simplified): For simplicity, since kernels that are multiples of $(1,2,2)$ are commonly used, we use abbreviated notation such as k1 that only denotes the temporal dimension.
-Specifically, k1 represents k1h2w2 (the kernel $(1,2,2)$), k2 represents k2h4w4 (the kernel $(2,4,4)$), k4 represents k4h8w8 (the kernel $(4,8,8)$), etc.
+Specifically, k1 represents k1h2w2 (the kernel $(1,2,2)$), k2 represents k2h4w4 (the kernel $(2,4,4)$), etc.
 Encoding frames: The notation f16k4 indicates that $16$ frames are encoded by the kernel k4 (the simplified k4h8w8) with kernel size $(4,8,8)$.
-Packing: The notation f16k4f4k2f1k1 shows a way to encode $19$ contiguous frames with three compression levels: the first $16$ frames are encoded by the kernel k4 (the kernel $(4,8,8)$), the next $4$ frames are encoded by the kernel k2 (the kernel $(2,4,4)$), and the last $1$ frame is encoded by the kernel k1 (the kernel $(1,2,2)$).
+Packing: The notation f16k4f2k2f1k1 shows a way to encode $19$ contiguous frames: the first $16$ frames are encoded by the kernel k4 (the kernel $(4,8,8)$), the next $2$ frames are encoded by the kernel k2 (the kernel $(2,4,4)$), and the last $1$ frame is encoded by the kernel k1 (the kernel $(1,2,2)$).
 Tail: We append the notation with td, ta, or tc to indicate the tail frames before or after packing, such as td_f16k4f4k2f1k1.
@@
 Skipping: The notation x skips an arbitrary number of frames (including 0 frames).
+History Discretization: The notation +d means converting history into discrete space.
 Generating: The notation g9 means generating 9 frames.
-With the above naming convention, we can represent all ablative structures in a compact form.
-Note that this naming also implies the sampling approach as discussed in Section 3.3.
-Consider the three similar structures and their sampling:
+With the above naming convention, we can represent all ablative structures in a compact form. Note that this naming also implies the sampling approach as discussed in Section 4.1:
 td_f16k4f4k2f1k1_g9: The vanilla sampling that generates frames in temporal order.
+td_f16k4f4k2f1k1_g9+D: The vanilla sampling with history discretization.
 td_f16k4f4k2f1k1_g9_x_f1k1: The anti-drifting sampling with an endpoint frame.
@@
 We implement both the text-to-video and image-to-video structures, though both are naturally supported by next-frame-section prediction models and do not need architecture modifications.
-The Wan base model is the official Wan2.1 model.
-The HunyuanVideo model is an improved version of official HunyuanVideo models to match the capability of Wan2.1, with the modifications: (1) adding the SigLip-Vision model google/siglip-so400m-patch14-384 as a vision encoder, (2) removing the reliance on Tencent’s internal MLLM, (3) freezing LLama3.1 as a pure text model and ignoring its multi-modality, and (4) continued training on high-quality data.
-Note that the numerical results presented in this paper should not be interpreted as direct comparisons between Wan and HunyuanVideo, as both were trained with similar computational budgets while Wan’s larger model size means fewer relative resources (e.g., batch size).
-Both models demonstrate comparable quality after sufficient training.
-We recommend HunyuanVideo as the default configuration for more efficient training and faster inference.
-Dataset
-We follow the guidelines of LTXVideo [13]’s dataset collection pipeline to gather data at multiple resolutions and quality levels.
-All data are filtered using quality measurements and motion scores to ensure a high-quality diverse distribution.
-We use aspect ratio bucketing for multi-resolution training with a minimum unit size of $32$ pixels.
-For example, the buckets at 480p resolution include: $(416,960)$, $(448,864)$, $(480,832)$, $(512,768)$, $(544,704)$, $(576,672)$, $(608,640)$, $(640,608)$, $(672,576)$, $(704,544)$, $(768,512)$, $(832,480)$, $(864,448)$, and $(960,416)$.
-Small-scale Training
-FramePack achieves a batch size of 64 on a single 8×A100-80G node with the 13B HunyuanVideo model at 480p resolution (without using any image-based workaround training that many community LoRAs are built on top of).
-This batch size is comparable even to image diffusion models like the 12B Flux, making FramePack suitable for personal or laboratory-scale training and experimentation.
-Training
-We conduct all experiments using H100 GPU clusters.
-Each FramePack variant in the ablation studies consumes approximately 48 hours of training, while the final models are trained for one week.
-We use the Adafactor optimizer with a fixed learning rate of 1e-5 without learning rate scheduling.
-The gradient norm clip is set to 0.5.
-Lower flow shift
-Since the next-frame-section prediction methods generate smaller 3D tensors at each inference compared to full-video diffusion, the models can be trained with much lower flow shift values.
-We use Flux’s dynamic flow shift to train all models.
-We find that this leads to sharp and clean results that are closer to real videos.
-4.3 Evaluation Metrics
+We report results with HunyuanVideo in the main paper (see also supplementary for Wan results).
+We conduct all experiments using H100 GPU clusters with training details in the supplementary.
+Note that FramePack achieves a batch size of about 64 on a single 8×A100-80G node with the 13B HunyuanVideo model at 480p resolution LoRA training with window size 2 or 3 (or batch size 32 of window size 4 or 5), making FramePack suitable for personal or laboratory-scale training and experimentation.
+We follow the guidelines of LTXVideo [17]’s dataset collection pipeline to gather data at multiple resolutions and quality levels (see also supplementary for more details).
+5.3 Quantitative Evaluation
 We discuss the metrics for evaluating ablative architectures.
@@
 For quantitative tests, we by default use 30 seconds for long videos and 5 seconds for short videos.
-4.3.1 Multi-dimension Metrics
-Multiple metrics for video evaluations are consistent with common benchmarks, e.g., VBench [18], VBench2 [61], etc. All scores are normalized using the same numerical system with VBench.
-Clarity: The MUSIQ [22] image quality predictor trained on the SPAQ [10] dataset. This metric measures artifacts such as noise and blurring in the generated frames.
-Aesthetic: The LAION aesthetic predictor [33]. This metric measures the aesthetic values perceived with a CLIP-based estimator.
+Metrics
+Multiple metrics for video evaluations are consistent with common benchmarks, e.g., VBench [24], VBench2 [75], etc.
+Clarity: The MUSIQ [27] image quality predictor trained on SPAQ [14].
+This metric measures artifacts like noise and blurring.
+Aesthetic: The LAION aesthetic predictor [40].
+This metric measures the aesthetic values perceived by a CLIP-based estimator.
 Motion: The video frame interpolation model [23] modified by VBench to measure the smoothness of motion.
@@
 Note that the “dynamic” metric and “motion” metric represent a trade-off, e.g., a still image may rank high on motion smoothness but will be penalized by low dynamic degrees.
-Semantic: The video-text score computed by ViCLIP [42]. This metric measures the overall semantic consistency between the generated video and the input text.
+Semantic: The video-text score computed by ViCLIP [53].
+This metric measures the overall semantic consistency between the generated video and the prompt.
 Anatomy: The ViT [13] pretrained by VBench for identifying the per-frame presence of hands, faces, bodies, etc.
-Identity: The facial feature similarity using ArcFace [7] with face detection by RetinaFace [8] to measure the identity consistency.
-Table 1: Ablation study.
-We evaluate different FramePack configurations across multiple global metrics, drifting metrics, and human assessments.
-The table is divided into three groups based on sampling approach: vanilla sampling (top), anti-drifting sampling (middle), and inverted anti-drifting sampling (bottom).
-The tests are conducted with HunyuanVideo as base.
-Bests in bold.
-ELO differences within ±$\pm 16$ are considered ties.
-Variant Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ td_f8k8f4k4f2k2f1k1_g9 67.33% 65.94% 94.53% 92.91% 20.06% 66.97% 71.72% 3.25% 3.45% 7.45% 16.56% 1084 8 td_f64k8f16k4f4k2f1k1_g9 67.71% 66.71% 95.09% 93.91% 21.39% 66.56% 71.50% 3.22% 3.38% 7.25% 16.43% 1068 9 td_f512k8f64k4f8k2f1k1_g9 67.74% 66.44% 94.27% 92.91% 21.90% 67.73% 71.35% 3.18% 3.32% 7.05% 15.95% 1088 8 td_f512k16f64k8f8k2f1k1_g9 66.13% 64.52% 94.77% 94.18% 19.21% 65.72% 71.51% 3.25% 3.45% 7.85% 17.62% 1072 9 td_f16k4f4k2f1k1_g9 67.37% 65.05% 95.97% 91.66% 19.15% 65.38% 71.73% 3.22% 3.48% 6.65% 17.36% 1076 9 td_f16k4f2k2f1k1_g1 65.81% 64.39% 94.91% 94.92% 19.20% 64.16% 69.70% 3.43% 3.77% 9.81% 20.89% 1036 11 td_f16k4f2k2f1k1_g4 66.57% 64.99% 94.00% 82.85% 19.40% 65.97% 69.06% 3.36% 3.68% 8.55% 19.09% 1052 10 td_f16k4f2k2f1k1_g9 67.15% 65.29% 94.08% 92.97% 20.73% 66.46% 71.08% 3.18% 3.42% 7.45% 18.05% 1080 9 tc_f16k4f2k2f1k1_g9 67.62% 65.07% 94.02% 90.33% 21.71% 71.70% 71.01% 3.15% 3.25% 7.25% 17.21% 1092 8 ta_f16k4f2k2f1k1_g9 67.02% 66.44% 94.11% 91.09% 21.48% 71.92% 72.18% 3.12% 3.18% 7.05% 17.66% 1096 8 td_f8k8f4k4f2k2f1k1_g9_x_f1k1 68.46% 66.95% 96.46% 75.55% 22.88% 85.10% 75.84% 2.95% 2.85% 5.95% 15.87% 1132 4 td_f64k8f16k4f4k2f1k1_g9_x_f1k1 68.21% 67.75% 95.74% 85.97% 22.79% 81.09% 76.29% 2.92% 2.98% 5.95% 15.99% 1136 4 td_f512k8f64k4f8k2f1k1_g9_x_f1k1 68.62% 67.72% 95.05% 76.50% 22.44% 82.01% 76.65% 2.88% 2.92% 5.75% 15.94% 1116 5 td_f512k16f64k8f8k2f1k1_g9_x_f1k1 68.35% 67.89% 97.73% 76.48% 22.75% 78.40% 76.16% 2.85% 2.85% 5.55% 14.04% 1120 5 td_f16k4f4k2f1k1_g9_x_f1k1 68.42% 67.59% 96.74% 74.37% 23.69% 79.14% 77.37% 2.82% 2.78% 5.35% 14.90% 1140 4 td_f16k4f2k2f1k1_g1_x_f1k1 65.32% 67.97% 95.26% 81.69% 19.97% 74.57% 77.93% 2.78% 2.72% 4.95% 14.80% 1084 7 td_f16k4f2k2f1k1_g4_x_f1k1 69.92% 67.49% 97.84% 71.90% 21.12% 74.84% 77.53% 2.75% 2.65% 4.95% 14.98% 1100 6 td_f16k4f2k2f1k1_g9_x_f1k1 69.51% 69.15% 96.97% 77.41% 23.03% 83.10% 69.25% 2.72% 2.58% 4.75% 13.73% 1144 4 tc_f16k4f2k2f1k1_g9_x_f1k1 69.62% 68.42% 96.45% 82.27% 23.08% 81.68% 69.08% 2.68% 2.52% 4.55% 13.54% 1148 4 ta_f16k4f2k2f1k1_g9_x_f1k1 69.21% 68.84% 97.87% 76.22% 22.77% 81.70% 75.23% 2.65% 2.45% 4.35% 13.76% 1152 4 f1k1_x_g9_f1k1f2k2f4k4f8k8_td 69.62% 67.87% 97.93% 88.79% 23.73% 86.99% 78.63% 2.55% 2.35% 4.15% 12.71% 1205 1 f1k1_x_g9_f1k1f4k2f16k4f64k8_td 69.56% 67.48% 97.42% 86.68% 24.48% 86.35% 78.06% 2.45% 2.25% 3.95% 12.52% 1209 1 f1k1_x_g9_f1k1f8k2f64k4f512k8_td 69.35% 67.89% 97.85% 88.21% 24.66% 76.99% 79.56% 2.35% 2.15% 3.75% 12.13% 1213 1 f1k1_x_g9_f1k1f8k2f64k8f512k16_td 69.20% 68.76% 99.11% 89.18% 24.35% 77.62% 79.88% 2.35% 2.05% 3.55% 11.10% 1239 1 f1k1_x_g9_f1k1f4k2f16k4_td 69.25% 67.40% 98.59% 89.01% 24.24% 77.31% 79.16% 2.45% 1.95% 3.35% 9.22% 1221 1 f1k1_x_g1_f1k1f2k2f16k4_td 69.74% 67.04% 98.09% 79.85% 24.89% 77.27% 80.86% 2.55% 2.15% 3.75% 11.37% 1148 3 f1k1_x_g4_f1k1f2k2f16k4_td 70.28% 68.11% 98.30% 79.45% 24.87% 77.95% 80.23% 2.45% 2.05% 3.75% 11.12% 1164 2 f1k1_x_g9_f1k1f2k2f16k4_td 70.73% 67.97% 98.11% 88.76% 25.79% 78.45% 84.01% 2.35% 1.95% 3.65% 11.98% 1225 1 f1k1_x_g9_f1k1f2k2f16k4_tc 70.48% 68.74% 98.16% 89.18% 27.01% 87.21% 81.04% 2.18% 1.85% 2.54% 11.71% 1229 1 f1k1_x_g9_f1k1f2k2f16k4_ta 70.81% 67.31% 98.15% 80.72% 25.60% 85.60% 82.76% 2.25% 1.77% 2.95% 9.85% 1233 1
-4.3.2 Drifting Measurements
-We discuss the methods to measure drifting.
-Several metrics from VBench-Long [19] already integrate long-range consistency measurements and may indicate drifting.
-We propose a more direct metric, namely start-end contrast, to measure drifting from multiple aspects.
+Identity: The facial feature similarity using ArcFace [11] with face detection by RetinaFace [12].
+Drifting measurement
 We observe that when drifting occurs, a significant difference emerges between the beginning and ending portions of a video across various quality metrics.
@@
 Since video models may generate frames in different temporal orders (either forward or backward), we use the absolute difference to ensure our metric remains direction-agnostic.
-4.3.3 Human Assessments
+Human assessments
 We collect human preferences from A/B tests.
@@
 We report ELO-K32 score and the relative ranking.
-4.4 Ablative Results
-As shown in Table 1, we note several discoveries: (1) The inverted anti-drifting sampling method achieves the best results in 5 out of 7 metrics, while other sampling methods achieve at most a single best metric.
-(2) The inverted anti-drifting sampling achieves the best performance in all drifting metrics.
-(3) Human evaluations indicate that generating 9 frames per section yields better perception than generating 1 or 4 frames, as evidenced by the higher ELO scores for configurations with g9 compared to their g1 or g4 counterparts.
-(4) While vanilla sampling achieved the highest dynamic score, this is likely attributable to drifting effects rather than genuine quality.
-(5) We also observe that differences between specific configuration options within the same sampling approach are relatively small and random, suggesting that the sampling category contributes more to the overall performance difference.
+Ablative results
+As shown in Table 1, we note several discoveries.
+(1) The inverted anti-drifting sampling method achieves the best results in 4 out of 7 metrics, and achieves the best performance in all drifting metrics.
+(2) However, the inverted anti-drifting sampling has a relatively small dynamic range.
+(3) While vanilla sampling achieved the highest dynamic score, this is likely attributable to drifting effects rather than genuine quality, evidenced by relatively low ELO scores.
+(4) The vanilla sampling with discrete history achieves highly competitive human scores while having a much larger dynamic range.
+(5) We also observe that differences between specific configuration options within the same sampling approach are relatively small and random, suggesting that the overall architecture contributes more to the general difference.
+History discretization parameter
+The history discretization is influenced by the parameter $K$ with higher $K$ giving stronger anti-drifting effects but also more challenging learning for smooth transitions between sections.
+In our tests, $K=128$ gives strong drift reduction with relatively minimal training difficulties.
+We provide more detailed ablations for $K$ in the supplementary materials.
+Additional evaluations
+We test feature-similarity-based packing with video game (world model) benchmarks in the supplementary material. See also supplementary material for Wan results and more details of decomposed metrics.
+Table 1: Ablation study.
+We evaluate different FramePack configurations across multiple global metrics, drifting metrics, and human assessments.
+The table is divided into 4 groups based on sampling approach: vanilla sampling, anti-drifting sampling, and inverted anti-drifting sampling, and vanilla sampling with discrete history.
+The tests are conducted with HunyuanVideo as base.
+Bests in bold.
+ELO differences within $\pm 16$ are considered ties.
+Variant Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank* $\downarrow$ td_f8k8f4k4f2k2f1k1_g9 67.33% 65.94% 94.53% 92.91% 20.06% 66.97% 71.72% 3.25% 3.45% 7.45% 16.56% 1090 5 td_f64k8f16k4f4k2f1k1_g9 67.71% 66.71% 95.09% 93.91% 21.39% 66.56% 71.50% 3.22% 3.38% 7.25% 16.43% 1070 5 td_f512k8f64k4f8k2f1k1_g9 67.74% 66.44% 94.27% 92.91% 21.90% 67.73% 71.35% 3.18% 3.32% 7.05% 15.95% 1085 5 td_f512k16f64k8f8k2f1k1_g9 66.13% 64.52% 94.77% 94.18% 19.21% 65.72% 71.51% 3.25% 3.45% 7.85% 17.62% 1068 5 td_f16k4f4k2f1k1_g9 67.37% 65.05% 95.97% 91.66% 19.15% 65.38% 71.73% 3.22% 3.48% 6.65% 17.36% 1072 5 td_f16k4f2k2f1k1_g1 65.81% 64.39% 94.91% 94.92% 19.20% 64.16% 69.70% 3.43% 3.77% 9.81% 20.89% 1030 6 td_f16k4f2k2f1k1_g4 66.57% 64.99% 94.00% 82.85% 19.40% 65.97% 69.06% 3.36% 3.68% 8.55% 19.09% 1050 6 td_f16k4f2k2f1k1_g9 67.15% 65.29% 94.08% 92.97% 20.73% 66.46% 71.08% 3.18% 3.42% 7.45% 18.05% 1074 5 tc_f16k4f2k2f1k1_g9 67.62% 65.07% 94.02% 90.33% 21.71% 71.70% 71.01% 3.15% 3.25% 7.25% 17.21% 1088 5 ta_f16k4f2k2f1k1_g9 67.02% 66.44% 94.11% 91.09% 21.48% 71.92% 72.18% 3.12% 3.18% 7.05% 17.66% 1092 5 td_f8k8f4k4f2k2f1k1_g9_x_f1k1 68.46% 66.95% 96.46% 75.55% 22.88% 85.10% 75.84% 2.95% 2.85% 5.95% 15.87% 1135 3 td_f64k8f16k4f4k2f1k1_g9_x_f1k1 68.21% 67.75% 95.74% 85.97% 22.79% 81.09% 76.29% 2.92% 2.98% 5.95% 15.99% 1140 3 td_f512k8f64k4f8k2f1k1_g9_x_f1k1 68.62% 67.72% 95.05% 76.50% 22.44% 82.01% 76.65% 2.88% 2.92% 5.75% 15.94% 1118 4 td_f512k16f64k8f8k2f1k1_g9_x_f1k1 68.35% 67.89% 97.73% 76.48% 22.75% 78.40% 76.16% 2.85% 2.85% 5.55% 14.04% 1115 4 td_f16k4f4k2f1k1_g9_x_f1k1 68.42% 67.59% 96.74% 74.37% 23.69% 79.14% 77.37% 2.82% 2.78% 5.35% 14.90% 1138 3 td_f16k4f2k2f1k1_g1_x_f1k1 65.32% 67.97% 95.26% 81.69% 19.97% 74.57% 77.93% 2.78% 2.72% 4.95% 14.80% 1080 5 td_f16k4f2k2f1k1_g4_x_f1k1 69.92% 67.49% 97.84% 71.90% 21.12% 74.84% 77.53% 2.75% 2.65% 4.95% 14.98% 1100 4 td_f16k4f2k2f1k1_g9_x_f1k1 69.51% 69.15% 96.97% 77.41% 23.03% 83.10% 69.25% 2.72% 2.58% 4.75% 13.73% 1142 3 tc_f16k4f2k2f1k1_g9_x_f1k1 69.62% 68.42% 96.45% 82.27% 23.08% 81.68% 69.08% 2.68% 2.52% 4.55% 13.54% 1145 3 ta_f16k4f2k2f1k1_g9_x_f1k1 69.21% 68.84% 97.87% 76.22% 22.77% 81.70% 75.23% 2.65% 2.45% 4.35% 13.76% 1150 3 f1k1_x_g9_f1k1f2k2f4k4f8k8_td 69.62% 67.87% 97.93% 88.79% 23.73% 86.99% 78.63% 2.55% 2.35% 4.15% 12.71% 1210 1 f1k1_x_g9_f1k1f4k2f16k4f64k8_td 69.56% 67.48% 97.42% 86.68% 24.48% 86.35% 78.06% 2.45% 2.25% 3.95% 12.52% 1215 1 f1k1_x_g9_f1k1f8k2f64k4f512k8_td 69.35% 67.89% 97.85% 88.21% 24.66% 76.99% 79.56% 2.35% 2.15% 3.75% 12.13% 1220 1 f1k1_x_g9_f1k1f8k2f64k8f512k16_td 69.20% 68.76% 99.11% 89.18% 24.35% 77.62% 79.88% 2.35% 2.05% 3.55% 11.10% 1235 1 f1k1_x_g9_f1k1f4k2f16k4_td 69.25% 67.40% 98.59% 89.01% 24.24% 77.31% 79.16% 2.45% 1.95% 3.35% 9.22% 1225 1 f1k1_x_g1_f1k1f2k2f16k4_td 69.74% 67.04% 98.09% 79.85% 24.89% 77.27% 80.86% 2.55% 2.15% 3.75% 11.37% 1150 3 f1k1_x_g4_f1k1f2k2f16k4_td 70.28% 68.11% 98.30% 79.45% 24.87% 77.95% 80.23% 2.45% 2.05% 3.75% 11.12% 1175 2 f1k1_x_g9_f1k1f2k2f16k4_td 70.73% 67.97% 98.11% 88.76% 25.79% 78.45% 84.01% 2.35% 1.95% 3.65% 11.98% 1228 1 f1k1_x_g9_f1k1f2k2f16k4_tc 70.48% 68.74% 98.16% 89.18% 27.01% 87.21% 81.04% 2.18% 1.85% 2.54% 11.71% 1230 1 f1k1_x_g9_f1k1f2k2f16k4_ta 70.81% 67.31% 98.15% 80.72% 25.60% 85.60% 82.76% 2.25% 1.77% 2.95% 9.85% 1232 1 td_f8k8f4k4f2k2f1k1_g9+D 69.12% 66.53% 96.91% 93.07% 24.79% 82.58% 73.34% 2.43% 2.74% 4.45% 14.37% 1225 1 td_f64k8f16k4f4k2f1k1_g9+D 69.35% 65.99% 97.21% 92.83% 22.43% 81.88% 72.90% 2.22% 2.44% 4.49% 13.74% 1170 2 td_f512k8f64k4f8k2f1k1_g9+D 69.26% 67.24% 96.28% 91.01% 24.63% 82.58% 71.07% 2.26% 2.01% 5.46% 13.00% 1169 2 td_f512k16f64k8f8k2f1k1_g9+D 69.93% 67.26% 97.60% 93.61% 23.87% 81.27% 73.39% 2.93% 2.92% 5.11% 15.70% 1139 3 td_f16k4f4k2f1k1_g9+D 69.49% 67.25% 96.95% 92.39% 23.40% 88.67% 72.08% 2.30% 2.63% 6.74% 13.51% 1223 1 td_f16k4f2k2f1k1_g1+D 68.16% 65.47% 95.94% 93.15% 23.97% 75.24% 71.78% 2.66% 3.70% 6.40% 16.83% 1145 3 td_f16k4f2k2f1k1_g4+D 68.16% 65.80% 95.26% 92.97% 22.97% 69.87% 71.91% 2.70% 3.73% 5.69% 17.06% 1142 3 td_f16k4f2k2f1k1_g9+D 69.78% 67.22% 96.91% 91.39% 24.40% 77.62% 74.95% 2.30% 2.49% 4.12% 14.11% 1222 1 tc_f16k4f2k2f1k1_g9+D 69.42% 66.24% 96.01% 92.53% 23.06% 79.15% 72.76% 2.87% 2.43% 5.51% 13.67% 1218 1 ta_f16k4f2k2f1k1_g9+D 68.51% 67.77% 96.90% 93.92% 23.24% 79.90% 70.01% 2.27% 2.41% 4.31% 15.85% 1216 1 * Based on the ELO scores and tie rules, the ranks are divided into: 1030-1050 (Rank 6), 1068-1092 (Rank 5), 1100-1118 (Rank 4), 1135-1150 (Rank 3), 1169-1175 (Rank 2), and 1210-1235 (Rank 1).
 5.4 Comparison to Alternative Architectures
+Table 2: Comparison with relevant methods.
+We compare across the same global metrics, drifting metrics, and human assessments.
+The tests are conducted with HunyuanVideo as base.
+Bests in bold.
+ELO differences within $\pm 16$ are considered ties.
+Method Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ Repeating image-to-video 56.73% 56.15% 94.34% 91.21% 17.74% 69.41% 73.06% 9.51% 3.92% 9.95% 19.88% 1015 5 Anchor frames (resembling StreamingT2V [20]) 69.58% 67.35% 99.96% 74.97% 25.76% 85.01% 79.52% 2.85% 2.15% 3.45% 9.25% 1173 2 Causal attention (resembling CausVid [65]) 62.88% 59.41% 96.98% 88.27% 19.15% 72.74% 75.32% 7.45% 3.15% 6.75% 15.96% 1087 4 DiffusionForcing [6] ($\sigma_{\text{train}}$ random, $\sigma_{\text{test}}=0.1$) 66.08% 65.76% 96.32% 91.59% 23.14% 75.93% 74.47% 4.84% 2.54% 3.33% 10.99% 1170 2 DiffusionForcing [6] ($\sigma_{\text{train}}$ random, $\sigma_{\text{test}}=0.5$) 67.41% 66.66% 92.93% 91.08% 24.03% 77.83% 76.42% 3.55% 2.39% 3.48% 9.40% 1174 2 DiffusionForcing [6] ($\sigma_{\text{train}}$ random, $\sigma_{\text{test}}=0$) 66.99% 64.73% 96.47% 90.45% 21.09% 80.62% 78.77% 8.41% 3.80% 8.47% 17.44% 1095 4 DiffusionForcing [6] ($\sigma_{\text{train}}=0.1$, $\sigma_{\text{test}}=0.1$) 66.19% 68.60% 94.89% 91.89% 22.49% 76.12% 78.27% 6.82% 3.79% 5.08% 10.78% 1149 3 History guidance (resembling HistoryGuidance [41]) 68.05% 68.74% 97.01% 73.39% 24.88% 81.84% 83.42% 7.35% 2.21% 5.25% 12.78% 1152 3 Inverted anti-drifting (f1k1_x_g9_f1k1f2k2f16k4) 71.15% 68.71% 99.45% 89.29% 28.15% 86.53% 82.11% 2.25% 1.85% 2.68% 8.58% 1220 1 Vanilla + discrete history (f16k4f2k2f1k1_g9+D, $K=256$) 70.01% 68.76% 95.65% 91.74% 28.37% 86.41% 82.22% 3.13% 2.05% 2.89% 8.74% 1224 1
 We discuss several relevant alternatives to generate videos in various ways.
@@
 Causal attention: Finetune full attention into causal attention for easier KV cache and faster inference. We implement a structure that resembles CausVid [65].
-Noisy history: Delay the denoising timestep on history latents so that history latents are noisy (but less noisy than the current generating latents).
-Reducing the reliance on the history is beneficial for interrupting error accumulation, thus mitigates drifting, but at the cost of aggravating forgetting.
-We implement a structure that resembles DiffusionForcing [4].
+DiffusionForcing: We conduct a detailed ablation study with DiffusionForcing [6].
+We focus on the history noise scheduling using the same scheduling as SkyreelV2 [7] that multiplies the diffusion timestep on history frames with $\sigma_{\text{train}}$ during training and $\sigma_{\text{test}}$ in inference to delay the denoising on history latents.
+Intuitively, $\sigma_{\text{test}}=0$ is equivalent to clean latents for history (no noise added to the history).
+We consider these ablations: (1) $\sigma_{\text{train}}$ being random with $\sigma_{\text{test}}=0.1$; (2) $\sigma_{\text{train}}$ being random with $\sigma_{\text{test}}=0.5$; (3) $\sigma_{\text{train}}$ being random with $\sigma_{\text{test}}=0$; (4) $\sigma_{\text{train}}=0.1$ and $\sigma_{\text{test}}=0.1$.
+Usually higher $\sigma_{\text{test}}$ reduces the reliance on the history, which is beneficial for interrupting error accumulation, thus mitigates drifting, but at the cost of aggravating forgetting.
 History guidance: Delay the denoising timestep on history latents but also put the completely noised history on the unconditional side of CFG guidance.
-This will speed up error accumulation thus aggravating drifting, but also enhance memory to mitigate forgetting.
+This will speed up error accumulation, thus aggravating drifting, but also enhance memory to mitigate forgetting.
 We implement a structure that resembles HistoryGuidance [41].
-Table 2: Comparison with relevant architectures.
-We compare with other relevant methods across the same global metrics, drifting metrics, and human assessments used in the ablation study.
-The tests are conducted with HunyuanVideo as base.
-Bests in bold.
-ELO differences within ±$\pm 16$ are considered ties.
-Method Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ Repeating image-to-video 56.73% 56.15% 94.34% 65.56% 17.74% 69.41% 73.06% 9.51% 3.92% 9.95% 19.88% 1016 5 Anchor frames (resembling StreamingT2V [15]) 69.58% 67.35% 99.96% 74.97% 25.76% 85.01% 79.52% 2.85% 2.15% 3.45% 9.25% 1172 2 Causal attention (resembling CausVid [53]) 62.88% 59.41% 96.98% 67.56% 19.15% 72.74% 75.32% 7.45% 3.15% 6.75% 15.96% 1087 4 Noisy history (resembling DiffusionForcing [4]) 66.75% 63.78% 97.20% 91.45% 22.98% 77.60% 77.92% 4.35% 2.05% 3.15% 9.85% 1146 3 History guidance (resembling HistoryGuidance [34]) 68.05% 68.74% 97.01% 73.39% 24.88% 81.84% 83.42% 7.35% 2.21% 5.25% 12.78% 1151 3 Proposed (f1k1_x_g9_f1k1f2k2f16k4) 71.15% 68.71% 99.45% 89.29% 28.15% 86.53% 82.11% 2.25% 1.85% 2.68% 8.58% 1221 1
-As shown in Table 2, we observe several findings: (1) The proposed method achieves the best results in 3 global metrics, while other methods excel in at most one or two metrics.
-(2) The proposed method achieves the best results across all drifting metrics.
-(3) This evaluation aligns with human perceptions as evidenced by the ELO score.
+As shown in Table 2, we observe several findings.
+(1) The inverted anti-drifting sampling achieves the best results across all drifting metrics, while having a relatively small dynamic range.
+(2) The vanilla sampling with discrete history is very competitive in drifting measurements, while having a relatively larger dynamic range.
+(3) Human perception prefers the two proposed candidates as evidenced by the ELO score.
+(4) See also the supplementary material for more detailed explanations, analysis, and comparisons with DiffusionForcing candidates.
 6 Conclusion
 In this paper, we presented FramePack, a neural network structure that aims to address the forgetting-drifting dilemma in next-frame prediction models for video generation.
-FramePack applies progressive compression to input frames based on their importance, ensuring the total context length converges to a fixed upper bound regardless of video duration.
-This is achieved through manipulating transformer patchify kernel sizes for different compression rates.
-Combined with the anti-drifting sampling methods that incorporate bi-directional context through early-established endpoints or inverted temporal ordering, our approach enables longer video generation while maintaining unchanged computational bottlenecks.
-Experiments suggest that FramePack can improve model responsiveness and visual quality in inference, while allowing for higher batch sizes in training.
+FramePack applies progressive compression to input frames based on their importance, ensuring the context length converges to a fixed upper bound.
+We discussed both time-proximity-based packing and feature-similarity-based packing.
+We also discussed the anti-drifting training methods with history discretization, and the anti-drifting sampling methods using bi-directional context planning and scheduling.
+Experiments suggest that FramePack can process a large number of frames, improve model responsiveness, and allow for higher batch sizes in training.
 The approach is compatible with existing video diffusion models and supports various compression variants that can be optimized for wider applications.
+Acknowledgments and Disclosure of Funding
+This work was partially supported by the Brown Institute for Media Innovation, by Google through their affiliation with Stanford Institute for Human-centered Artificial Intelligence (HAI) and by a Hoffman-Yee HAI grant.